# Zero-Shot Floor-Aware Object-Goal Navigation via LLM-Driven Coarse-to-Fine Exploration

### v1

@@ 1 Introduction
To quantify the importance of multi-floor navigation, we analyze the validation sets of two widely used embodied navigation benchmarks: HM3D [37] and MP3D [9]. We find that 65.0% of validation episodes in HM3D and 54.7% in MP3D involve multi-level buildings, with 28.1% and 19.5%, respectively, requiring explicit floor transitions (see Appendix  C.1). These findings highlight the critical need for methods capable of effectively addressing multi-floor navigation in the OGN task. However, existing approaches often fail to meet this requirement.
Existing OGN methods can be broadly categorized into learning-based and zero-shot approaches, yet both fail to address zero-shot multi-floor navigation effectively. Specifically: (1) Learning-based methods [10, 56, 38, 40, 13, 39, 48] are typically trained on specific object categories and environments, which limits their ability to generalize to unseen settings. Our experiments reveal a significant performance drop when these models are applied to new environments (see Appendix B.2). (2) Zero-shot methods [31, 60, 64, 11, 59, 61, 57, 58] mostly overlook the possibility that target objects may reside on different floors or lack the necessary mechanisms for cross-floor navigation. Recent work such as MFNP [63] treated stairs as irreversible paths, preventing agents from revisiting previously explored floors even when semantic evidence suggests a higher likelihood of finding targets there. This restricts exploration scope and reduces efficiency in multi-floor scenarios.
@@
In addition, navigating in open-world environments introduces another challenge: ZS-OGN requires agents to interpret novel object descriptions using general knowledge. While Large Language Models (LLMs) have shown promise as planners for ZS-OGN, most LLM-based approaches rely on object detectors to convert 3D visual observations into textual descriptions before inputting them to the LLM planner. This process introduces modality discrepancies and leads to information loss. An alternative approach VLFM [59] use vision-language models for image-text matching at each step, selecting the frontier point most relevant to the target description. While this avoids explicit object detection, it introduces two key limitations. First, the local matching strategy may yield suboptimal global paths, causing the agent to oscillate between distant frontiers and reducing both Success Rate (SR) and Success weighted by Path Length (SPL). Second, decisions are made greedily based on current observations, lacking long-term planning and explainabilily.
@@
  • Our method achieves state-of-the-art performance for ZS-OGN on HM3D and MP3D benchmarks, improving SR by $7.1\%$ and SPL by $6.8\%$ on HM3D, and $3.4\%$ in SR and $0.1\%$ in SPL on MP3D. We further demonstrate its practicality through sim-to-real deployment on a quadruped robot.
@@ 2 Related Work
Object-Goal Navigation. This task involves locating and reaching a target object in unseen environments [6], requiring perception, spatial reasoning, and action planning [15, 19, 25, 17]. As a core capability of embodied AI, it supports downstream tasks such as vision-language navigation [4] and physical interaction [18]. Most prior work employs Reinforcement Learning (RL) [10, 54, 52] or Imitation Learning (IL) [40, 39, 48]. While effective in controlled settings, these methods face key limitations in real-world deployment: RL requires extensive reward annotations [56], IL depends on impractical expert demonstrations [49], and both assume a closed-world setting with fixed object categories [21], failing when encountering novel objects [22].
@@
LLMs for Navigation. Recent advances in Large Language Models (LLMs) enable strong zero-shot generalization in embodied navigation. L3MVN [60] introduces commonsense reasoning, while TriHelper [64] improves long-term planning with dynamic modules. VoroNav [50] combines geometric planning with LLM-based scoring. Recently, GAMap [61] uses affordance attributes for goal localization, abd Recently, SG-Nav [57] proposes a 3D scene graph-based prompting strategy for LLM reasoning. However, current LLM-based methods face two main limitations: (1) they primarily operate on single-floor maps without explicit inter-floor modeling, restricting vertical navigation; and (2) frequent LLM calls to evaluate all frontiers result in high computational cost and limited scalability. In contrast, our Coarse-to-Fine Frontier Reasoning module uses a semantic-value map to filter promising frontiers and invokes the LLM only when necessary to improve efficiency.
@@ 3 Methodology
This section firstly introduces the OGN task and presents the overall architecture of ASCENT. We then describe our online Multi-Floor Spatial Abstraction approach for multi-floor planning, concluding with a Coarse-to-Fine Frontier Reasoning strategy that leverages exploration value maps and contextual understanding for zero-shot navigation.
@@ 3.3.2 Multi-Floor Topology Modeling
Multi-Floor Map Manager. This module maintains dynamic BEV map lists $\mathcal{L}_{\mathrm{obs}}=[\mathcal{M}^{(1)}_{\mathrm{obs}},...,\mathcal{M}^{(N)}_{\mathrm{obs}}]$ and $\mathcal{L}_{\mathrm{val}}=[\mathcal{M}^{(1)}_{\mathrm{val}},...,\mathcal{M}^{(N)}_{\mathrm{val}}]$, where $N$ denotes the total number of explored floors and each entry corresponds to a floor-specific BEV map. In contrast to prior work MFNP [63], which might merged multi-floor data into a single representation, leading to spatial overlap or semantic confusion. Instead, our design preserves per-floor geometry and semantic history to ensure consistent spatial mapping across different floors.
@@ 4.1 Experimental Settings
Implementation Details. Our experiments are conducted on Habitat 3.0 [35], building upon the VLFM architecture [59]. We use D-FINE [34] for COCO-class objects [26] and Grounding-DINO [28] for open-vocabulary detection. Object contours are segmented by Mobile-SAM [62], and stair detection integrates a fine-tuned RedNet [20] with RGB-D inputs. BLIP-2 [24] computes semantic similarity, and Qwen2.5-7B-Instruct [55] performs LLM-based reasoning. All experiments run on a single compute node with two NVIDIA RTX 3090 GPUs.
@@ 4.2 Comparisons to State-of-the-Arts
Although our performance lags behind leading supervised methods such as XGX [48] and RIM [13], our framework operates entirely without task-specific training data, that makes it well-suited for sim-to-real deployment. As shown in Appendix B.2, these supervised approaches exhibit significant performance drops when transferred to unseen environments or novel object categories, whereas our method remains robust across domains. This confirms that ASCENT has achieved strong performance without labeled data and demonstrated superior generalization capabilities, setting a new standard in zero-shot, floor-aware navigation.
@@ 4.3 Qualitative Analysis
Figure 4: Comparisons of methods in multi-floor environments. (a) VLFM [59] fails to navigate to another floor, resulting in a low success rate (SR). (b) Adding Multi-Floor Spatial Abstraction module to VLFM [59] allows reaching different floors but often leads to unnecessarily long paths. (c) Our method efficiently navigates across floors and selects the nearly optimal path to the target.
@@
4.4 Ablation Study
Effect of Both Intra-Floor and Inter-Floor Strategies. As shown in Tab. 4, our method achieves consistent improvements over the baseline approach VLFM and the state-of-the-art zero-shot method MFNP across all scenarios of single-floor and multi-floor settings. In single-floor settings, ASCENT attains $76.1\%$ SR ($6.3\%$ absolute gain vs VLFM) with $22\%$ closer to targets than VLFM, demonstrating enhanced intra-floor exploration through our Coarse-to-Fine Frontier Reasoning. The multi-floor results reveal even more significant advantages: ASCENT outperforms MFNP by $7.1\%$ SR and $7.4\%$ SPL while reducing navigation distance by $11.6\%$, validating our framework’s effectiveness in resolving inter-floor transitions.
@@ 4.4 Ablation Study
MFNP Qwen-VLChat-Int4 Qwen2-7B 58.3 26.7 41.1 15.4
MFNP† BLIP-2 Qwen2.5-7B 55.8 24.9 38.5 12.4
Ours BLIP-2 Qwen2.5-7B 65.4 33.5 44.5 15.5
@@
Effect of Architecture Design. Tab. 3 presents the ablation study of our framework components on the HM3D dataset. In addition to standard metrics, we report Distance to the Goal (DTG), which measures the geodesic distance from the final position to the closest target, offering finer insights into localization accuracy. Removing the Multi-Floor Topology Modeling (MFTM) module leads to the most significant performance drop: -8.7% SR and -4.9% SPL, highlighting its critical role in multi-floor navigation. Disabling Coarse-to-Fine Frontier Reasoning (CFFR) module results in a -7.7% SR and -5.0% SPL decline, underscoring its importance in guiding efficient exploration. Without the Exploration Cost Map (ECM), the agent suffers -9.1% SR and -5.8% SPL degradation, indicating its necessity in avoiding costly paths. When prior knowledge (PK) is removed, the performance decreases only moderately (-3.3% SR and -3.4% SPL). This suggests that while LLMs inherently encode sufficient world knowledge and spatial reasoning capabilities for zero-shot navigation, the integration of statistical priors provides a structured reference that further aligns the agent’s behavior with environmental regularities.
Effect of Large Models. We conduct an ablation study on the large model in Tab. 3, by replacing the original Instruction Interpolator components in MFNP (the previous SOTA method) with our model’s configuration: substituting Qwen-VLChat-Int4 with BLIP-2 for vision and Qwen2-7B with Qwen2.5-7B for language processing. Surprisingly, this modification (MFNP*) leads to performance degradation, with SR dropping from 58.3% to 55.8% and SPL decreasing from 26.7% to 24.9%.
This ablation highlights that our improvements stem from novel algorithm design rather than simply employing more powerful foundation models.
@@ A.2 Motivation of Prior Knowledge
A.3 More Details of Object-Goal Task
- MOVE_FORWARD: Move forward by $0.25\,\mathrm{m}$,
- TURN_LEFT: Rotate left by $30^{\circ}$,
- TURN_RIGHT: Rotate right by $30^{\circ}$,
- LOOK_UP: Tilt camera upward by $30^{\circ}$,
- LOOK_DOWN: Tilt camera downward by $30^{\circ}$,
- STOP: Terminate the episode.
@@ A.4 Vertical Localization Constraints
IMU-derived height estimates are prone to drift and lack precision during dynamic movements like stair climbing [7]. Visual SLAM also struggles in texture-poor environments such as staircases, where feature extraction becomes unreliable [32]. Even if vertical coordinates were available, determining floor levels would still require knowledge of building-specific floor heights. However, prior information typically unknown in unseen environments. Small sensor errors could lead to incorrect floor classification and planning failures.
@@ A.5 Implementation Details of Coarse-Grained Frontier Proposal
To further improve efficiency and reduce redundancy, we apply visual similarity filter to remove the frontiers whose representative images stay below a threshold $\theta$ on structural overlap. This design priorities frontiers based on scene semantics and spatial diversity:
$\mathcal{F}_{\mathrm{coarse}}=\left\{f\in\mathcal{F}_{\mathrm{boundary}}\,\middle|\,\text{SSIM}(f,\mathcal{F}_{\mathrm{cache}})<\tau_{ssim}\right\},$ (6)
@@ Appendix B Additional Ablation Experiments
Table 6: Comparison Results of Different Object Detectors.
@@
VLFM GroundingDINO + YOLOv7 $52.5$ $30.4$ $36.4$ $17.5$
@@
ASCENT(Ours) GroundingDINO + YOLOv7 $60.9$ $29.6$ $42.4$ $13.8$
@@
We also conduct an ablation study on the object detector in Tab. 6, revealing three key insights: (1) When using YOLOv7 [45] (equivalent to VLFM’s detector), our framework still outperforms VLFM by $8.4\%$ SR, validating our architecture’s superiority beyond sensing capabilities; (2) Adopting the state-of-the-art object detector D-FINE [34] further improves navigation performance (+$4.5\%$ SR), confirming our framework’s ability to leverage detection advances. (3) We employ a perfect object detector, i.e. ground truth semantic segmentation sensor, which results in an additional $4.9\%$ improvement in SR and a $7.6\%$ increase in SPL.
@@
To evaluate the generalization ability of closed-set methods and our method, we compare the most advanced learning-based methods (RIM[13], PIRLNav [39] and XGX [48]) with our zero-shot approach on both HM3D and MP3D datasets. As shown in Tab. 7, the learning-based methods exhibit strong performance on the training dataset (HM3D) but suffer significant drops in SR and SPL when transferred to MP3D, highlighting their struggle to generalize to unseen environments. In contrast, our zero-shot method maintains robust performance across both datasets, achieving the best SR on MP3D and demonstrating superior generalization ability.
@@ C.1 Motivation of Multi-Floor Navigation
Figure 6: Multi-Floor Scene and Episode Analysis: Over half of scenarios involve multiple floors, with approximately one-fifth of episodes requiring cross-floor capabilities for the agent, consistent across the training and validation sets of two public benchmarks.
@@
Figure 7: Capability comparision of navigation methods.

### v2

@@ 1 Introduction
To quantify the importance of multi-floor navigation, we analyze the validation sets of two widely used embodied navigation benchmarks: HM3D [38] and MP3D [9]. We find that $65.0\%$ of validation episodes in HM3D and $54.7\%$ in MP3D involve multi-level buildings, with $28.1\%$ and $19.5\%$, respectively, requiring explicit floor transitions (see Appendix  C.1). These findings highlight the critical need for methods capable of effectively addressing multi-floor navigation in the OGN task. However, existing approaches often fail to meet this requirement.
Existing OGN methods can be broadly categorized into learning-based and zero-shot approaches, yet both fail to address zero-shot multi-floor navigation effectively. Specifically: (1) Learning-based methods [10, 57, 39, 41, 13, 40, 49] are typically trained on specific object categories and environments, which limits their ability to generalize to unseen settings. Our experiments reveal a significant performance drop when these models are applied to new environments (see Appendix B.2). (2) Zero-shot methods [32, 61, 65, 11, 60, 62, 58, 59] mostly overlook the possibility that target objects may reside on different floors or lack the necessary mechanisms for cross-floor navigation. Recent work, such as MFNP [64], treated stairs as irreversible paths, preventing agents from revisiting previously explored floors even when semantic evidence suggests a higher likelihood of finding targets there. This restricts exploration scope and reduces efficiency in multi-floor scenarios.
@@
In addition, navigating in open-world environments introduces another challenge: ZS-OGN requires agents to interpret novel object descriptions using general knowledge. While Large Language Models (LLMs) have shown promise as planners for ZS-OGN, most LLM-based approaches rely on object detectors to convert 3D visual observations into textual descriptions before inputting them to the LLM planner. This process introduces modality discrepancies and leads to information loss. An alternative approach, VLFM [60], uses vision-language models for image-text matching at each step, selecting the frontier point most relevant to the target description. While this avoids explicit object detection, it introduces two key limitations. First, the local matching strategy may yield suboptimal global paths, causing the agent to oscillate between distant frontiers and reducing both Success Rate (SR) and Success Weighted by Path Length (SPL). Second, decisions are made greedily based on current observations, lacking long-term planning and explainability.
@@
  • Our approach achieves state-of-the-art performance for ZS-OGN on HM3D and MP3D benchmarks, improving SR by $7.1\%$ and SPL by $6.8\%$ on HM3D, and $3.4\%$ in SR and $0.1\%$ in SPL on MP3D. We further demonstrate its practicality through sim-to-real deployment on a quadruped robot.
@@ 2 Related Work
Object-Goal Navigation. This task involves locating and reaching a target object in unseen environments [6], requiring perception, spatial reasoning, and action planning [15, 19, 25, 17, 26]. As a core capability of embodied AI, it supports downstream tasks such as vision-language navigation [4] and physical interaction [18]. Most prior work employs Reinforcement Learning (RL) [10, 55, 53] or Imitation Learning (IL) [41, 40, 49]. While effective in controlled settings, these methods face key limitations in real-world deployment: RL requires extensive reward annotations [57], IL depends on impractical expert demonstrations [50], and both assume a closed-world setting with fixed object categories [21], failing when encountering novel objects [22].
@@
LLMs for Navigation. Recent advances in Large Language Models (LLMs) enable strong zero-shot generalization in embodied navigation. L3MVN [61] introduces commonsense reasoning, while TriHelper [65] improves long-term planning with dynamic modules. VoroNav [51] combines geometric planning with LLM-based scoring. Recently, GAMap [62] uses affordance attributes for goal localization. SG-Nav [58] proposes a 3D scene graph-based prompting strategy for LLM reasoning. However, current LLM-based methods face two main limitations: (1) they primarily operate on single-floor maps without explicit inter-floor modeling, restricting vertical navigation; and (2) frequent LLM calls to evaluate all frontiers result in high computational cost and limited scalability. In contrast, our Coarse-to-Fine Frontier Reasoning module uses a semantic-value map to filter promising frontiers and invokes the LLM only when necessary to improve efficiency.
@@ 3 Methodology
This section firstly introduces the OGN task and presents the overall architecture of ASCENT. We then describe our online Multi-Floor Spatial Abstraction approach for multi-floor planning, concluding with a Coarse-to-Fine Frontier Reasoning strategy that leverages exploration value maps and contextual understanding for zero-shot navigation. An overview of our proposed framework is shown in Fig. 2.
@@ 3.3.2 Multi-Floor Topology Modeling
Multi-Floor Map Manager. This module maintains dynamic BEV map lists $\mathcal{L}_{\mathrm{obs}}=[\mathcal{M}^{(1)}_{\mathrm{obs}},...,\mathcal{M}^{(N)}_{\mathrm{obs}}]$ and $\mathcal{L}_{\mathrm{val}}=[\mathcal{M}^{(1)}_{\mathrm{val}},...,\mathcal{M}^{(N)}_{\mathrm{val}}]$, where $N$ denotes the total number of explored floors and each entry corresponds to a floor-specific BEV map. In contrast to prior work, MFNP [64], might merge multi-floor data into a single representation, leading to spatial overlap or semantic confusion. Instead, our design preserves per-floor geometry and semantic history to ensure consistent spatial mapping across different floors.
@@ 4.1 Experimental Settings
Implementation Details. Our experiments are conducted on Habitat $3.0$ [36], building upon the VLFM architecture [60]. We use D-FINE [35] for COCO-class objects [27] and Grounding-DINO [29] for open-vocabulary detection. Object contours are segmented by Mobile-SAM [63], and stair detection integrates a fine-tuned RedNet [20] with RGB-D inputs. BLIP-$2$ [24] computes semantic similarity, and Qwen2.52.52.52.5-$7$B-Instruct [56] performs LLM-based reasoning. All experiments were run on a single compute node with two NVIDIA RTX $3090$ GPUs.
@@ 4.2 Comparisons to State-of-the-Arts
Although our performance lags behind leading supervised methods such as XGX [49] and RIM [13], our framework operates entirely without task-specific training data, which makes it well-suited for sim-to-real deployment. As shown in Appendix B.2, these supervised approaches exhibit significant performance drops when transferred to unseen environments or novel object categories, whereas our method remains robust across domains. This confirms that ASCENT has achieved strong performance without labeled data and demonstrated superior generalization capabilities, setting a new standard in zero-shot, floor-aware navigation.
@@ 4.3 Qualitative Analysis
Figure 4: Comparisons of methods in multi-floor environments. (a) VLFM [60] fails to navigate to another floor, resulting in a low success rate (SR). (b) Adding the Multi-Floor Spatial Abstraction module to VLFM [60] allows reaching different floors but often leads to unnecessarily long paths. (c) Our method efficiently navigates across floors and selects the nearly optimal path to the target.
@@
MFNP Qwen-VLChat-Int4444 Qwen2222-$7$B $58.3$ $26.7$ $41.1$ $15.4$
MFNP† BLIP-$2$ Qwen2.52.52.52.5-$7$B $55.8$ $24.9$ $38.5$ $12.4$
Ours BLIP-$2$ Qwen2.52.52.52.5-$7$B $65.4$ $33.5$ $44.5$ $15.5$
4.4 Ablation Study
Effect of Both Intra-Floor and Inter-Floor Strategies. As shown in Tab. 4, our method achieves consistent improvements over the baseline approach VLFM and the state-of-the-art zero-shot method MFNP, across all scenarios of single-floor and multi-floor settings. In single-floor settings, ASCENT attains $76.1\%$ SR ($6.3\%$ absolute gain vs VLFM) with $22\%$ closer to targets than VLFM, demonstrating enhanced intra-floor exploration through our Coarse-to-Fine Frontier Reasoning. The multi-floor results reveal even more significant advantages: ASCENT outperforms MFNP by $7.1\%$ SR and $7.4\%$ SPL while reducing navigation distance by $11.6\%$, validating the effectiveness of our framework in resolving inter-floor transitions.
@@ 4.4 Ablation Study
Effect of Architecture Design. Tab. 3 presents the ablation study of our framework components on the HM3D dataset. In addition to standard metrics, we report Distance to the Goal (DTG), which measures the geodesic distance from the final position to the closest target, offering finer insights into localization accuracy. Removing the Multi-Floor Topology Modeling (MFTM) module leads to the most significant performance drop: -$8.7\%$ SR and -$4.9\%$ SPL, highlighting its critical role in multi-floor navigation. Disabling the Coarse-to-Fine Frontier Reasoning (CFFR) module results in a -$7.7\%$ SR and -$5.0\%$ SPL decline, underscoring its importance in guiding efficient exploration.
Without the Exploration Cost Map (ECM), the agent suffers -$9.1\%$ SR and -$5.8\%$ SPL degradation, indicating its necessity in avoiding costly paths. When prior knowledge (PK) is removed, the performance decreases only moderately (-$3.3\%$ SR and -$3.4\%$ SPL). This suggests that while LLMs inherently encode sufficient world knowledge and spatial reasoning capabilities for zero-shot navigation, the integration of statistical priors provides a structured reference that further aligns the agent’s behavior with environmental regularities.
Effect of Large Models. We conduct an ablation study on the large model in Tab. 3, by replacing the original Instruction Interpolator components in MFNP (the previous SOTA method) with the configuration of our model: substituting Qwen-VLChat-Int4444 with BLIP-$2$ for vision and Qwen2222-$7$B with Qwen2.52.52.52.5-$7$B for language processing. Surprisingly, this modification (MFNP*) leads to performance degradation, with SR dropping from $58.3\%$ to $55.8\%$ and SPL decreasing from $26.7\%$ to $24.9\%$. This ablation highlights that our improvements stem from novel algorithm design rather than simply employing more powerful foundation models.
@@ A.2 Motivation of Prior Knowledge
A.3 More Details of Object-Goal Task
  • MOVE-FORWARD: Move forward by $0.25\,\mathrm{m}$.
  • TURN-LEFT: Rotate left by $30^{\circ}$.
  • TURN-RIGHT: Rotate right by $30^{\circ}$.
  • LOOK-UP: Tilt camera upward by $30^{\circ}$.
  • LOOK-DOWN: Tilt camera downward by $30^{\circ}$.
  • STOP: Terminate the episode.
@@ A.4 Vertical Localization Constraints
IMU-derived height estimates are prone to drift and lack precision during dynamic movements like stair climbing [7]. Visual SLAM also struggles in texture-poor environments such as staircases, where feature extraction becomes unreliable [33]. Even if vertical coordinates were available, determining floor levels would still require knowledge of building-specific floor heights. However, prior information is typically unknown in unseen environments. Small sensor errors could lead to incorrect floor classification and planning failures.
@@ A.5 Implementation Details of Coarse-Grained Frontier Proposal
To further improve efficiency and reduce redundancy, we apply a visual similarity filter to remove the frontiers whose representative images stay below a threshold $\theta$ on structural overlap. This design prioritizes frontiers based on scene semantics and spatial diversity:
$\mathcal{F}_{\mathrm{coarse}}=\left\{f\in\mathcal{F}_{\mathrm{boundary}}\,\middle|\,\text{SSIM}(f,\mathcal{F}_{\mathrm{cache}})<\tau_{\mathrm{ssim}}\right\},$ (6)
@@ Appendix B Additional Ablation Experiments
Table 6: Comparative results of different object detectors.
@@
VLFM GroundingDINO + YOLOv7777 $52.5$ $30.4$ $36.4$ $17.5$
@@
ASCENT (Ours) GroundingDINO + YOLOv7777 $60.9$ $29.6$ $42.4$ $13.8$
@@
We also conduct an ablation study on the object detector in Tab. 6, revealing three key insights: (1) When using YOLOv7 [46] (equivalent to VLFM’s detector), our framework still outperforms VLFM by $8.4\%$ SR, validating our architecture’s superiority beyond sensing capabilities; (2) Adopting the state-of-the-art object detector D-FINE [35] further improves navigation performance (+$4.5\%$ SR), confirming our framework’s ability to leverage detection advances. (3) We employ a perfect object detector, i.e., ground truth semantic segmentation sensor, which results in an additional $4.9\%$ improvement in SR and a $7.6\%$ increase in SPL.
@@
To evaluate the generalization ability of closed-set methods and our method, we compare the most advanced learning-based methods (RIM[13], PIRLNav [40], and XGX [49]) with our zero-shot approach on both HM3D and MP3D datasets. As shown in Tab. 7, the learning-based methods exhibit strong performance on the training dataset (HM3D) but suffer significant drops in SR and SPL when transferred to MP3D, highlighting their struggle to generalize to unseen environments. In contrast, our zero-shot method maintains robust performance across both datasets, achieving the best SR on MP3D and demonstrating superior generalization ability.
@@ C.1 Motivation of Multi-Floor Navigation
Figure 6: Multi-Floor Scene and Episode Analysis. We observe that over half of the scenarios involve multiple floors, with approximately one-fifth of episodes requiring cross-floor capabilities for the agent, consistent across the training and validation sets of two public benchmarks.
@@
Figure 7: Capability comparisons of navigation methods.